OpenStudy (anonymous):

A particle is moving on an axis Ox. From time t=0 to time t=32 the particle is travelling at a constant speed 15ms-1. The particle then decelerates from 15ms-1 to rest in T seconds Find the value of T @Engineering

OpenStudy (asnaseer):

I don't think you have given enough information to solve this problem. e.g. if it decelerated at a very slow rate of deceleration, then it would take it a lot longer to come to rest than if it decelerated at a very high rate of deceleration.

OpenStudy (anonymous):

Assuming acceleration is constant... u = 15m/s, v = 0m/s, t = T. v = u + at 0 = 15 + at at = -15 t = -15/a = T As asnaseer said, you need some more information to get any more than an expression.